Recovery in progress at the beginning of the shift. From 21:29 UTC to 21:58 UTC A. Bertolini remotely recovered the vertical position of SDB2 that was decreasing due to the intervention in the afternoon.
After this activity we tried to relock again but unfortunately the IMC was very unstable, LOW_NOISE_1 reached again 23:37 UTC.
The ITF systematically unlocked while locking the first OMC. J. Casanueva investigated the alignment of the ITF but no major issue were found. Under request of the Run Coordinator I contacted the DET Oncall ( R. Bonnard ) at 2:35 UTC and R. Gouaty at 4:07 UTC.
Scan of the OMC in progress.
WE LC opened at 23:42 UTC after unlock. Loops closed at 23:48 UTC.
22:19 UTC - 22:53 UTC
0:37 UTC - 1:15 UTC
2:28 UTC - 3:00 UTC
4:10 UTC - 4:40 UTC
DMS: Tank / VAC..WE1_LN2_BAR: to be monitored. If the value increased up to 2.5 contact VAC OnCall.